# Learning Encoding-Decoding Direction Pairs to Unveil Concepts of Influence in Deep Vision Networks

### v2

@@ A.13 Hyper-Parameter Study with Respect to Target Separation Margin
Tables 17 and 18 consider varying the target margin $\tau^{mm}$ when learning EDDP. For EfficientNet, increasing the target classification margin (i.e. the linear separability of patch embeddings) leads to a substantial improvement in both $\mathcal{S}^{1}$ and $\mathcal{S}^{2}$ for both EDDP-C and EDDP-U. We stress the fact that increasing the margin is accomplished by decreasing $\tau^{mm}$, as the margin is inversely proportional to the target loss. For ResNet18, the results are more mixed: for EDDP-U, decreasing the margin improves $\mathcal{S}^{1}$ and $\mathcal{S}^{2}$, while for EDDP-C there is an improvement in $\mathcal{S}^{1}$ but at the cost of a slightly inferior $\mathcal{S}^{2}$ score. We emphasize that when $\tau^{mm}=4.0$, the results for EfficientNet are substantially better than the metrics we reported in the main body of the paper, for which we used $\tau^{mm}=5.0$ (Section A.12).
A.14 Ablation Study with Respect to Filter-Signal Orthogonality Loss
@@
A.15 Comparison with Unsupervised Interpretable Basis Extraction and Concept-Basis Extraction in Practical Experiments
Table 25: Comparison and Ablation with respect to method enhancements that we propose in this work. $\mathcal{L}^{al}$ stands for Augmented Lagrangian Loss and $\mathcal{L}^{cc}$ refers to the CNN Classifier loss that was proposed in CBE. Methods below the dashed line correspond to variations that use the contributions we make in this work. Here the network is ResNet18 and $I=512$.
@@
Table 26: Comparison and Ablation with respect to method enhancements that we propose in this work. $\mathcal{L}^{al}$ stands for Augmented Lagrangian Loss and $\mathcal{L}^{cc}$ refers to the CNN Classifier loss that was proposed in CBE. Methods below the dashed line correspond to variations that use the contributions we make in this work. Here the network is ResNet50 and $I=2048$.
@@
Tables 25 and 26 briefly present comparisons of our work with the previous works that we extend, in terms of interpretability. First, with other aspects of Concept Basis Extraction (CBE) (CBE) being intact, we assess the efficacy of our Unconstrained Uncertainty Region Alignment loss in improving the interpretability of the clustering, compared to the CNN Classifier Loss $\mathcal{L}^{cc}$ that was proposed previously in CBE. This case is referred to in the tables as CBE /w $\mathcal{L}^{uur}$. We observe that in three out of the four cases, our Uncertainty Region Alignment loss leads to a notable relative improvement, by up to 22.56% in $\mathcal{S}^{2}$, which indicates a clustering with improved concept diversity. When we additionally take into account the rest of our contributions, and compare EDDP-U with CBE, we find that in the same three out of four cases, interpretability is further improved by up to 58.8% in terms of $\mathcal{S}^{1}$. Transitioning from CBE to EDDP-C, may lead to further improvements in $\mathcal{S}^{2}$ by up to 33.5%, yet with a decreased score in terms of $\mathcal{S}^{1}$.
A.16 More Qualitative Segmentations and Statistics for Evaluating the Interpretability of the Concept Detectors
Figures 29, 30, 31, 32, 33, 34, 37, 38, 39 depict qualitative segmentation results obtained using the concept detectors that were learned with the proposed method for the architectures that we studied in this work. Visualizations are obtained using NetworkDissection. We can verify that concept detectors appear to be monosemantic. In some cases, there are more than one concept detector detecting the same concept. However, in most cases, the sets of positively classified samples across detectors are disjoint. Figures 35 and 36 depict qualitative segmentations obtained via NMF and PCA. In some of the depicted examples, it is evident that the identified concepts are less monosemantic. Tables 27, 28, 29 and 30 summarize the statistics obtained by Network Dissection regarding the interpretability of the clusterings considered in this work.
@@ II Method Color Object Part Material Action Scene Texture Total
A.17 More Global Model Explanations via Concept Sensitivity Testing
@@
A.18 More Local Explanations with Concept Contribution Maps
@@
A.19 Details on Counterfactual Explanation Example
@@
A.20 Details on Toy Model Correction
@@
A.21 Details on Computational Resources
@@
A.22 Relation to Sparse Auto-Encoders
The superposition hypothesis Superposition assumes that neural networks linearly represent more features than there are neurons in their hidden layers. In this line of work, a feature is defined to be an abstract property of the input, which may or may not align with human intuition, and is exploited by the network to make predictions. Sparse AutoEncoders (SAEs) DictionaryNLP have been proposed as a tool to take features out of superposition; that is, given the activations of a network’s intermediate layer, SAEs try to linearly decompose these activations in terms of latent feature components under a sparsity objective.
Let ${\bm{x}}\in{\mathbb{R}}^{D}$ denote the activation of a network in a hidden layer of study and $I\gg D$ denote the number of latent features that the network is assumed to represent. While several SAE variants have been proposed, such as SAE1; SAE2; SAE3; matryoshka-saes, in their baseline form DictionaryNLP they first extract latent feature components $\bm{v}$} by:

### v3

@@ A.14 Hyper-Parameter Study with Respect to Target Separation Margin
Tables 17 and 18 consider varying the target margin $\tau^{mm}$ when learning EDDP. For EfficientNet, increasing the target classification margin (i.e., the linear separability of patch embeddings) leads to substantial improvements in both $\mathcal{S}^{1}$ and $\mathcal{S}^{2}$ for both EDDP-C and EDDP-U. We stress that increasing the margin is accomplished by decreasing $\tau^{mm}$, as the margin is inversely proportional to the target loss. For ResNet18, the results are more mixed: for either EDDP-U or EDDP-C, decreasing the margin mostly improves $\mathcal{S}^{1}$ while keeping similar $\mathcal{S}^{2}$. We emphasize that when $\tau^{mm}=4.0$, the results for EfficientNet are substantially better than the metrics we reported in the main body of the paper, for which we used $\tau^{mm}=5.0$ (Section A.13).
A.15 Ablation Study with Respect to Filter-Signal Orthogonality Loss
@@
A.16 Comparison with Unsupervised Interpretable Basis Extraction and Concept-Basis Extraction in Practical Experiments
Table 25: Comparison and Ablation with respect to method enhancements that we propose in this work. $\mathcal{L}^{al}$ stands for Augmented Lagrangian Loss and $\mathcal{L}^{cc}$ refers to the CNN Classifier loss that was proposed in Doumanoglou et al. (2024). Methods below the dashed line correspond to variations that use the contributions we make in this work. Here the network is ResNet18 and $I=512$.
@@
Table 26: Comparison and Ablation with respect to method enhancements that we propose in this work. $\mathcal{L}^{al}$ stands for Augmented Lagrangian Loss and $\mathcal{L}^{cc}$ refers to the CNN Classifier loss that was proposed in Doumanoglou et al. (2024). Methods below the dashed line correspond to variations that use the contributions we make in this work. Here the network is ResNet50 and $I=2048$.
@@
Tables 25 and 26 briefly present comparisons of our work with the previous works that we extend, in terms of interpretability. First, with other aspects of Concept Basis Extraction (CBE) (Doumanoglou et al., 2024) being intact, we assess the efficacy of our Unconstrained Uncertainty Region Alignment loss in improving the interpretability of the clustering, compared to the CNN Classifier Loss $\mathcal{L}^{cc}$ that was proposed previously in Doumanoglou et al. (2024). This case is referred to in the tables as CBE /w $\mathcal{L}^{uur}$. We observe that in three out of the four cases, our Uncertainty Region Alignment loss leads to a notable relative improvement, by up to 22.56% in $\mathcal{S}^{2}$, which indicates a clustering with improved concept diversity. When we additionally take into account the rest of our contributions, and compare EDDP-U with CBE, we find that in the same three out of four cases, interpretability is further improved by up to 58.8% in terms of $\mathcal{S}^{1}$. Transitioning from CBE to EDDP-C, may lead to further improvements in $\mathcal{S}^{2}$ by up to 33.5%, yet with a decreased score in terms of $\mathcal{S}^{1}$.
A.17 More Qualitative Segmentations and Statistics for Evaluating the Interpretability of the Concept Detectors
Figures 29, 30, 31, 32, 33, 34, 37, 38, 39 depict qualitative segmentation results obtained using the concept detectors that were learned with the proposed method for the architectures that we studied in this work. Visualizations are obtained using Bau et al. (2017). We can verify that concept detectors appear to be monosemantic. In some cases, there are more than one concept detector detecting the same concept. However, in most cases, the sets of positively classified samples across detectors are disjoint. Figures 35 and 36 depict qualitative segmentations obtained via NMF and PCA. In some of the depicted examples, it is evident that the identified concepts are less monosemantic. Tables 27, 28, 29 and 30 summarize the statistics obtained by Network Dissection regarding the interpretability of the clusterings considered in this work.
@@ II Method Color Object Part Material Action Scene Texture Total
A.18 More Global Model Explanations via Concept Sensitivity Testing
@@
A.19 More Local Explanations with Concept Contribution Maps
@@
A.20 Details on Counterfactual Explanation Example
@@
A.21 Details on Toy Model Correction
@@
A.22 Details on Computational Resources
@@
A.23 Relation to Sparse Auto-Encoders
The superposition hypothesis Elhage et al. (2022) assumes that neural networks linearly represent more features than there are neurons in their hidden layers. In this line of work, a feature is defined to be an abstract property of the input, which may or may not align with human intuition, and is exploited by the network to make predictions. Sparse AutoEncoders (SAEs) Bricken et al. (2023) have been proposed as a tool to take features out of superposition; that is, given the activations of a network’s intermediate layer, SAEs try to linearly decompose these activations in terms of latent feature components under a sparsity objective.
Let ${\bm{x}}\in{\mathbb{R}}^{D}$ denote the activation of a network in a hidden layer of study and $I\gg D$ denote the number of latent features that the network is assumed to represent. While several SAE variants have been proposed, such as Lim et al. (2024); Cunningham et al. (2024); Sharkey et al. (2022); Bussmann et al. (2025), in their baseline form Bricken et al. (2023) they first extract latent feature components $\bm{v}$} by: